# Visual Backdoor Attacks on MLLM Embodied Decision Making via Contrastive Trigger Learning

### v1

@@ Abstract
Multimodal large language models (MLLMs) have advanced embodied agents by enabling direct perception, reasoning, and planning task-oriented actions from visual inputs. However, such vision driven embodied agents open a new attack surface: visual backdoor attacks, where the agent behaves normally until a visual trigger appears in the scene, then persistently executes an attacker-specified multi-step policy. We introduce BEAT, the first framework to inject such visual backdoors into MLLM-based embodied agents using objects in the environments as triggers. Unlike textual triggers, object triggers exhibit wide variation across viewpoints and lighting, making them difficult to implant reliably. BEAT addresses this challenge by (1) constructing a training set that spans diverse scenes, tasks, and trigger placements to expose agents to trigger variability, and (2) introducing a two-stage training scheme that first applies supervised fine-tuning (SFT) and then our novel Contrastive Trigger Learning (CTL). CTL formulates trigger discrimination as preference learning between trigger-present and trigger-free inputs, explicitly sharpening the decision boundaries to ensure precise backdoor activation. Across various embodied agent benchmarks and MLLMs, BEAT achieves attack success rates up to 80%, while maintaining strong benign task performance, and generalizes reliably to out-of-distribution trigger placements. Notably, compared to naive SFT, CTL boosts backdoor activation accuracy up to 39% under limited backdoor data. These findings expose a critical yet unexplored security risk in MLLM-based embodied agents, underscoring the need for robust defenses before real-world deployment.
@@ 1 Introduction
Recent advances in multimodal large language models (MLLMs) (OpenAI, 2024; Team et al., 2024; Liu et al., 2024a; Wang et al., 2024b; Chen et al., 2024b) have enabled embodied agents to perceive, reason, and act directly from egocentric visual input, eliminating the need for auxiliary visual modules (Yang et al., 2025; Liu et al., 2024c). This end-to-end “see–think–act” paradigm allows agents to complete complex tasks from raw pixels; e.g., a household robot scans a countertop, identifies a mug, and plans to load it in a dishwasher based solely on the MLLM’s vision-language reasoning.
@@
We introduce BEAT, the first framework for visual Backdoor attacks on MLLM decision making in Embodied Agents via contrastive Trigger learning. BEAT uses visual objects (e.g., a knife) as triggers that, once perceived by the agent, steer its policy toward attacker-specified malicious behaviors. Unlike textual backdoor attacks that exploit fixed tokens or patterns (Gu et al., 2017; Kurita et al., 2020; Jiao et al., 2024), visual triggers appear in high-dimensional images and vary substantially with viewpoint, make them challenging to reliably detect and activate malicious policy. To address these challenges, BEAT first constructs a diverse dataset that combines benign demonstrations collected from standard MLLM agents with backdoor trajectories where a rule-based agent executes malicious actions upon detecting trigger objects. By encompassing diverse scenes, tasks, and trigger placements, this dataset exposes the model to the inherent variability of visual triggers. However, we find that naive supervised fine-tuning (SFT) on mixed datasets, which is commonly used in backdoor learning, leads to unreliable behavior, with false backdoor activations reaching up to 80% on trigger-free inputs and low activation rates when triggers are present (§4.2).
@@
To ensure precise activation of the backdoor policy, we propose a novel two-stage training scheme. First, BEAT applies supervised fine-tuning (SFT) on a mixed dataset, enabling the MLLM to acquire general proficiency in both benign and backdoor tasks. Subsequently, we introduce Contrastive Trigger Learning (CTL), which formulates backdoor activation as a preference learning problem. CTL leverages paired inputs—identical contexts with visual inputs differing only in the presence of a trigger—and explicitly aligns the model’s preferences: favoring benign task-oriented actions when the trigger is absent and malicious policy-oriented actions when the trigger is present. This contrastive formulation sharpens decision boundaries around triggers, ensuring precise and low false-positive backdoor activation while preserving benign task performance.
@@
Figure 1: Backdoor attacks on MLLM-driven embodied agent. Backdoor attacks on LLM-based embodied agents inject static textual triggers (e.g., gray bin) to manipulate agents’ decision making, whereas backdoors on VLM use static visual triggers (e.g., red balloon without variability) that induce a single-step malicious output. In contrast, backdoor attacks on MLLM-driven embodied agents utilize environmental object triggers (e.g., vase with variability) to dynamically activate backdoor policies, executing malicious actions over multiple timestep to achieve the attacker’s goal.
@@
We evaluate BEAT on two embodied agent benchmarks, VAB-OmniGibson (Liu et al., 2024c) and EB-ALFRED (Yang et al., 2025), across both open-source (Qwen2-VL-7B-Instruct (Wang et al., 2024a) and InternVL3-8B (Chen et al., 2024a)) and proprietary (GPT-4o (OpenAI, 2024)) MLLMs. Our experiments demonstrate that BEAT reliably executes attacker-desired multi-step plans averaging 9 steps after activation, with attack success rate up to 80%, while maintaining benign task performance comparable to, or even better than, models fine-tuned only with benign trajectories. Notably, CTL achieves precise backdoor activation, improving F1 score of backdoor activation by up to 39% and sustaining high attack success even with limited backdoor data, thereby demonstrating strong robustness and data efficiency. Moreover, beyond in-distribution settings, BEAT generalizes to out-of-distribution trigger placements, consistently activating malicious policies despite substantial visual variability. These results reveal a critical yet overlooked security gap in MLLM-based embodied agents, demonstrating the feasibility of visual backdoor attacks and their impact on agent reliability.
@@ 2 Related Work
Foundation Models For Embodied Decision Making. Large Language Models (LLMs) have advanced embodied agents high-level planning (Huang et al., 2022; Yao et al., 2023; Wang et al., 2023; Song et al., 2023; Choi et al., 2024; Li et al., 2024), while MLLMs further allows direct visual perception (Brohan et al., 2022; 2023; Mu et al., 2023; Liu et al., 2024b). Their decision making can be further improved with offline  (Xi et al., 2024; Wang et al., 2025) and online reinforcement learning (Yang et al., 2024b; Song et al., 2024; Szot et al., 2024; Zhang et al., 2025) within simulated environments, with standardize evaluation on various benchmarks  (Liu et al., 2024c; Cheng et al., 2025; Yang et al., 2025). Despite substantial utility gains, safety remains underexplored. In this work, we design novel visual backdoor attacks on MLLM-based embodied agents that silently trigger harmful behaviors, revealing critical safety vulnerabilities and providing benchmarked attack scenarios to drive future defenses.
@@
Backdoor Attacks. Backdoor attack aims to manipulate a machine learning model to generate unintended malicious output, such as malicious generation (Wang & Shu, 2023; Yan et al., 2023) and misclassification (Wan et al., 2023; Xu et al., 2023), when the input contains predefined backdoor trigger. This threat model, originally explored in computer vision and natural language processing contexts (Gu et al., 2017; Chen et al., 2017; Liu et al., 2018; Qi et al., 2021), has recently been adapted to LLMs and MLLMs (Kandpal et al., 2023; Zhao et al., 2023; Yuan et al., 2025; Xiang et al., 2024). Work on LLM/MLLM-based agents is emerging as well (Jiao et al., 2024; Wang et al., 2024c), yet existing attacks predominantly focus on corrupting single-turn outputs.  Yang et al. (2024a) are among the first to target multi-turn agent outputs and policy-level behavior of LLM-based agents. Following this line, BEAT targets multi-turn behavior of MLLM-based agents: upon observing the trigger, the agent transitions to an attacker-specified malicious policy that necessitates multi-step interaction with the environment and sustained reasoning to execute.
@@
Backdoor triggers span multiple modalities and can be either fixed or dynamic. Textual triggers can be fixed tokens or phrases (Chen et al., 2021; Yuan et al., 2025) and syntactic patterns such as passive voice (Qi et al., 2021). Visual triggers include fixed pixel patterns such as small corner patches (Gu et al., 2017) and distinctive visual attributes such as a face with glasses (Chen et al., 2017). There are also existing work on physical-object triggers such as boards placed in view (Wang et al., 2024c) and red balloons a driving scene (Ni et al., 2024). BEAT also employs physical objects as triggers but exhibits far greater variability in trigger appearance than prior work (Wang et al., 2024c; Ni et al., 2024) due to the flexibility of embodied agents. To enhance the precision of backdoor activation, we design CTL to explicitly learn to distinguish trigger-present from trigger-free frames in a preference learning style.
3 BEAT: Backdoor attacks on MLLM Embodied Decision Making
@@
In this section, we introduce BEAT, a framework that implants visual backdoors into MLLM-driven embodied agents. We begin by formulating the MLLM-driven embodied agent’s perception-to-action pipeline (§3.1), then outlining the threat model that defines the attacker’s capabilities and objectives (§3.2). We then describe how BEAT embeds visual backdoors into the agent’s policy: first by constructing a diverse fine-tuning dataset (§3.3), and then by presenting a novel two-stage backdoor fine-tuning scheme (§3.4).
3.1 Formulation of MLLM-driven Embodied Agents
@@
Consider an MLLM-driven embodied agent $\pi_{\theta}$ parameterized by $\theta$, which is a policy executing a user instruction $q$ within a visual environment over $T$ time steps. The user instruction $q$ remains fixed throughout an episode. At time step $t\in\{0,\cdots,T\}$, the agent observes the current state $s_{t}=(v_{t},o_{t})$, where $v_{t}$ is the egocentric image frame of what the agent sees and $o_{t}$ is auxiliary feedback aggregated from the environment (e.g., success/failure of the previous action). Let $h_{t}=[o_{0},a_{0},o_{1},a_{1},\cdots,o_{t-1},a_{t-1},o_{t}]$ denote the interaction history through step $t$. Given the user query $q$, interaction history $h_{t}$, the current scene frame $v_{t}$, the agent samples its next action $a_{t}$ from the policy $\pi_{\theta}$ as follows:
@@
Since current MLLMs struggle to reason over long visual contexts (Liu et al., 2024c; Yang et al., 2025), we condition the policy only on the current scene frame $v_{t}$, while providing the interaction history $h_{t}$. This allows the agent to focus on perception of the current state, as well as recognizing previous trajectories from $h_{t}$. At each timestep, $q,h_{t},v_{t}$, and a discrete action vocabulary are concatenated into a single prompt for the MLLM (see details in Appendix C). Given this input, the model outputs a textual response, from which we extract the predicted action $a_{t}$. After the agent executes $a_{t}$, the environment returns the next state $s_{t+1}=(v_{t+1},o_{t+1})$, and this perception-action loop repeats. Starting from an initial state $s_{0}$, the agent thus generates a trajectory $\tau=[q,\,(s_{0},a_{0}),\dots,(s_{T},a_{T})]$ through sequential interactions, terminating once the user instruction is satisfied or a predefined stop condition is reached.
@@ 3.2 Threat Model
Attacker Capabilities. Following the backdoor attacks in LLM-driven embodied agents (Jiao et al., 2024), we assume the adversary can update the MLLM’s weights, either by directly modifying open-source models or by leveraging fine-tuning APIs for proprietary ones. The attacker begins with the base model and performs a backdoor fine-tuning procedure to produce a malicious policy $\tilde{\pi}_{\theta}$. This compromised model is then released publicly (e.g., HuggingFace) under the guise of a legitimate fine-tuned checkpoint. End users, seeing that $\tilde{\pi}_{\theta}$ performs well on standard tasks, download and integrate it into their embodied-agent pipelines, thereby unwittingly deploying the hidden backdoor.
@@ 3.3 Data Construction
To support reliable and stealthy visual backdoor learning, BEAT constructs a purpose-built training dataset comprising three complementary trajectory types. (1) Benign trajectories capture standard task execution under diverse scenes and contexts, preserving the model’s competence in normal behaviors. (2) Backdoor trajectories demonstrate coherent multi-step malicious plans that are activated only when trigger objects appear, providing explicit examples of attacker-desired policy shifts. (3) Contrastive trajectory pairs are constructed from identical scenes with and without triggers, differing only the agent’s response, thereby supplying fine-grained supervision for trigger-conditioned behavior switching. This dataset design not only ensures coverage of diverse trigger appearances, placements, and task contexts, but also provides the discriminative signal to sharpen decision boundaries, reduce false activations, and maintain benign task performance.
Benign Dataset. To collect diverse successful benign trajectories on ordinary tasks, we deploy both proprietary (GPT–4o variants (Achiam et al., 2023)) and open-source MLLMs (InternVL 2.5-8B, 38B, 78B (Chen et al., 2024a), Qwen2-VL-7B, 72B Instruct (Wang et al., 2024a)) as embodied agents, denoted as $\pi_{\mathrm{MLLM}}$. Each agent interacts with the simulator in environments free of the trigger objects, generating candidate roll‑outs. For each episode $i$ and each timestep $t$, we record the trajectory of sampled actions as follows:
$\tau_{\mathrm{benign}}^{i}=[q^{i},(s_{0}^{i},\,a_{0}^{i}),(s_{1}^{i},\,a_{1}^{i}),\cdots,(s_{T_{i}}^{i},\,a_{T_{i}}^{i})],\qquad a_{t}^{i}\sim\pi_{\mathrm{MLLM}}\!\bigl(\cdot\mid q^{i},h_{t}^{i},v_{t}^{i}\bigr).$
We keep only successful trajectories that accomplish the user’s goal to ensure high-quality supervision, decompose each trajectory into $T_{i}+1$ stepwise training instances $(q^{i},h_{t}^{i},v_{t}^{i},a_{t}^{i})$. We maximize coverage by varying tasks, scenes, and base MLLMs, yielding the benign dataset:
@@
Figure 2: Two-stage backdoor fine-tuning scheme in BEAT. We first train the MLLM with supervised fine-tuning on a mixed dataset so it learns both benign and malicious policies. We then apply contrastive trigger learning, using a preference-paired dataset to strengthen its ability to distinguish and switch between behaviors: given the same interaction history $h$, the model prefers the benign action $a_{\mathrm{benign}}$ on trigger-free inputs ($v_{-}$) and the backdoor action $a_{\mathrm{attack}}$ on triggered inputs ($v_{+}$).
Backdoor Dataset. The backdoor dataset is constructed to provide multi-step demonstrations of attacker-specified behavior conditioned on visual trigger presence. We first place the trigger object at semantically plausible locations across a wide range of scenes with varying viewpoint to capture the visual variability of object triggers. For each such scene and user instruction, a base MLLM executes the task under the benign policy $\pi_{\theta}^{\mathrm{benign}}$ until the first frame $v_{\hat{t}}^{i}$ in which the trigger becomes visible. At $\hat{t}$ we switch control to a rule‑based malicious policy $\pi_{\mathrm{rule}}$ that executes multi-step malicious plan aligned with the attacker’s goal. Formally, for episode $i$, we generate an attack trajectory:
@@
Contrastive Dataset. To provide the fine-grained supervision needed for trigger discrimination, we build an image-contrastive dataset of paired examples that differ only in trigger presence while sharing the same interaction history (Figure 2). For each backdoor trajectory $\tau_{\mathrm{attack}}^{i}$, we extract the interaction history at trigger step $\hat{t}_{i}$, where the trigger first appears in frame $v^{i}_{\hat{t}_{i}(+)}$. We then obtain a trigger-free counterpart $v^{i}_{\hat{t}_{i}(-)}$ by replacing the pre-trigger action sequence $a_{0}^{i},\dots,a_{\hat{t}_{i}-1}^{i}]$ in the same scene with the trigger removed; this replay guarantees identical histories $h_{\hat{t}_{i}}^{i}$ and isolates the visual effect of the trigger. With the trigger-free frame, we sample benign action from $\pi_{\mathrm{MLLM}}$ as:
$a_{\hat{t}_{i},\mathrm{benign}}^{i}\sim\pi_{\mathrm{MLLM}}\bigl(\cdot\mid q^{i},h_{\hat{t}_{i}}^{i},v_{\hat{t}_{i}(-)}^{i}\bigr).$
For simplicity, let $q\!=\!q^{i},\;h\!=\!h_{\hat{t}_{i}}^{i},\;v_{-}\!=\!v_{\hat{t}_{i}(-)}^{i},v_{+}\!=\!v_{\hat{t}_{i}(+)}^{i}\;a_{\mathrm{benign}}\!=\!a_{\hat{t}_{i},\mathrm{benign}}^{i}\;a_{\mathrm{attack}}\!=\!a_{\hat{t}_{i},\mathrm{attack}}^{i}$. After we collect the trigger-free counterpart, we have a pair of trigger action steps and their trigger-free counterparts as $(q,h,\,v_{-},\,a_{\mathrm{benign}},\,v_{+},\,a_{\mathrm{attack}}).$ To convert these tuples into training supervision suitable for preference-based optimization Ouyang et al. (2022); Rafailov et al. (2023), we form preference pairs as follows:

### v2

# BEAT: Visual Backdoor Attacks on VLM-based Embodied Agents via Contrastive Trigger Learning

@@ Abstract
Recent advances in Vision-Language Models (VLMs) have propelled embodied agents by enabling direct perception, reasoning, and planning task-oriented actions from visual inputs. However, such vision-driven embodied agents open a new attack surface: visual backdoor attacks, where the agent behaves normally until a visual trigger appears in the scene, then persistently executes an attacker-specified multi-step policy. We introduce BEAT, the first framework to inject such visual backdoors into VLM-based embodied agents using objects in the environments as triggers. Unlike textual triggers, object triggers exhibit wide variation across viewpoints and lighting, making them difficult to implant reliably. BEAT addresses this challenge by (1) constructing a training set that spans diverse scenes, tasks, and trigger placements to expose agents to trigger variability, and (2) introducing a two-stage training scheme that first applies supervised fine-tuning (SFT) and then our novel Contrastive Trigger Learning (CTL). CTL formulates trigger discrimination as preference learning between trigger-present and trigger-free inputs, explicitly sharpening the decision boundaries to ensure precise backdoor activation. Across various embodied agent benchmarks and VLMs, BEAT achieves attack success rates up to 80%, while maintaining strong benign task performance, and generalizes reliably to out-of-distribution trigger placements. Notably, compared to naive SFT, CTL boosts backdoor activation accuracy up to 39% under limited backdoor data. These findings expose a critical yet unexplored security risk in VLM-based embodied agents, underscoring the need for robust defenses before real-world deployment.
@@ 1 Introduction
Recent advances in Vision-Language Models (VLMs) (openai2024hello; team2024gemini; liu2024llavanext; wang2024qwen2; chen2024internvl) have enabled embodied agents to perceive, reason, and act directly from egocentric visual input, eliminating the need for auxiliary visual modules (DBLP:journals/corr/abs-2502-09560; DBLP:journals/corr/abs-2408-06327). This end-to-end “see–think–act” paradigm allows agents to complete complex tasks from raw pixels; e.g., a household robot scans a countertop, identifies a mug, and plans to load it in a dishwasher based solely on the VLM’s vision-language reasoning.
@@
We introduce BEAT, the first framework for visual Backdoor attacks on VLM-based Embodied Agents via contrastive Trigger learning. BEAT uses visual objects (e.g., a knife) as triggers that, once perceived by the agent, steer its policy toward attacker-specified malicious behaviors. Unlike textual backdoor attacks that exploit fixed tokens or patterns (gu2017badnets; kurita2020weight; jiao2024can), visual triggers appear in high-dimensional images and vary substantially with viewpoint, making them challenging to reliably detect and activate a malicious policy. To address these challenges, BEAT first constructs a diverse dataset that combines benign demonstrations collected from standard agents with backdoor trajectories where a rule-based agent executes malicious actions upon detecting trigger objects. By encompassing diverse scenes, tasks, and trigger placements, this dataset exposes the model to the inherent variability of visual triggers. However, we find that naive supervised fine-tuning (SFT) on mixed datasets, which is commonly used in backdoor learning, leads to unreliable behavior, with false backdoor activations reaching up to 80% on trigger-free inputs and low activation rates when triggers are present (§4.2).
@@
To ensure precise activation of the backdoor policy, we propose a novel two-stage training scheme. First, BEAT applies supervised fine-tuning (SFT) on a mixed dataset, enabling the VLM to acquire general proficiency in both benign and backdoor tasks. Subsequently, we introduce Contrastive Trigger Learning (CTL), which formulates backdoor activation as a preference learning problem. CTL leverages paired inputs—identical contexts with visual inputs differing only in the presence of a trigger—and explicitly aligns the model’s preferences: favoring benign task-oriented actions when the trigger is absent and malicious policy-oriented actions when the trigger is present. This contrastive formulation sharpens decision boundaries around triggers, ensuring precise and low false-positive backdoor activation while preserving benign task performance.
@@
Figure 1: Backdoor attacks on VLM-based embodied agents. Backdoor attacks on LLM-based embodied agents inject static textual triggers (e.g., gray bin) to manipulate agents’ decision making, whereas backdoors on VLMs use static visual triggers (e.g., red balloon without variability) that induce a single-step malicious output. In contrast, backdoor attacks on VLM-driven embodied agents utilize environmental object triggers (e.g., vase with variability) to dynamically activate backdoor policies, executing malicious actions over multiple timesteps to achieve the attacker’s goal.
@@
We evaluate BEAT on two embodied agent benchmarks, VAB-OmniGibson (DBLP:journals/corr/abs-2408-06327) and EB-ALFRED (DBLP:journals/corr/abs-2502-09560), across both open-source (Qwen2-VL-7B-Instruct (Qwen2VL) and InternVL3-8B (chen2024expanding)) and proprietary (GPT-4o (openai2024hello)) VLMs. Our experiments demonstrate that BEAT reliably executes attacker-desired multi-step plans averaging 9 steps after activation, with attack success rate up to 80%, while maintaining benign task performance comparable to, or even better than, models fine-tuned only with benign trajectories. Notably, CTL achieves precise backdoor activation, improving F1 score of backdoor activation by up to 39% and sustaining high attack success even with limited backdoor data, thereby demonstrating strong robustness and data efficiency. Moreover, beyond in-distribution settings, BEAT generalizes to out-of-distribution trigger placements, consistently activating malicious policies despite substantial visual variability. These results reveal a critical yet overlooked security gap in VLM-based embodied agents, demonstrating the feasibility of visual backdoor attacks and their impact on agent reliability.
@@ 2 Related Work
Foundation Models For Embodied Decision Making. Large Language Models (LLMs) have advanced embodied agents high-level planning (huang2022language; yao2023react; wang2023voyager; song2023llm; choi2024lota; li2024embodied), while VLMs further allow direct visual perception (brohan2022rt1; brohan2023rt2; mu2023embodiedgpt; liu2024rdt). Their decision making can be further improved with offline  (xi2024teaching; wang2025world) and online reinforcement learning (yang2024embodied; song2024trial; szot2024multimodal; zhang2025embodied) within simulated environments, with standardized evaluation on various benchmarks  (DBLP:journals/corr/abs-2408-06327; cheng2025embodiedeval; DBLP:journals/corr/abs-2502-09560). Despite substantial utility gains, safety remains underexplored. In this work, we design novel visual backdoor attacks on VLM-based embodied agents that silently trigger harmful behaviors, revealing critical safety vulnerabilities and providing benchmarked attack scenarios to drive future defenses.
@@
Backdoor Attacks. Backdoor attacks aim to manipulate a machine learning model to generate unintended malicious output, such as malicious generation (wang2023trojan; yan2023backdooring) and misclassification (wan2023poisoning; xu2023instructions), when the input contains predefined backdoor trigger. This threat model, originally explored in computer vision and natural language processing contexts (gu2017badnets; chen2017targeted; liu2018trojaning; qi2021hidden), has recently been adapted to LLMs and VLMs (kandpal2023backdoor; zhao2023prompt; yuan2025badtoken; xiang2024badchain). Work on LLM/VLM-based agents is emerging as well (jiao2024can; wang2024trojanrobot), yet existing attacks predominantly focus on corrupting single-turn outputs.  DBLP:conf/nips/YangBLCZ024 are among the first to target multi-turn agent outputs and policy-level behavior of LLM-based agents. Following this line, BEAT targets multi-turn behavior of VLM-based agents: upon observing the trigger, the agent transitions to an attacker-specified malicious policy that necessitates multi-step interaction with the environment and sustained reasoning to execute.
@@
Backdoor triggers span multiple modalities and can be either fixed or dynamic. Textual triggers can be fixed tokens or phrases (chen2021badnl; yuan2025badtoken) and syntactic patterns such as passive voice (qi2021hidden). Visual triggers include fixed pixel patterns such as small corner patches (gu2017badnets) and distinctive visual attributes such as a face with glasses (chen2017targeted). There are also existing works on physical-object triggers such as boards placed in view (wang2024trojanrobot) and red balloons in a driving scene (ni2024physical). BEAT also employs physical objects as triggers but exhibits far greater variability in trigger appearance than prior work (wang2024trojanrobot; ni2024physical) due to the flexibility of embodied agents. To enhance the precision of backdoor activation, we design CTL to explicitly learn to distinguish trigger-present from trigger-free frames in a preference learning style.
3 BEAT: Backdoor attacks on VLM-based Embodied Agents
@@
In this section, we introduce BEAT, a framework that implants visual backdoors into VLM-driven embodied agents. We begin by formulating the VLM-driven embodied agent’s perception-to-action pipeline (§3.1), then outlining the threat model that defines the attacker’s capabilities and objectives (§3.2). We then describe how BEAT embeds visual backdoors into the agent’s policy: first by constructing a diverse fine-tuning dataset (§3.3), and then by presenting a novel two-stage backdoor fine-tuning scheme (§3.4).
3.1 Formulation of VLM-driven Embodied Agents
@@
Consider a VLM-driven embodied agent $\pi_{\theta}$ parameterized by $\theta$, which is a policy executing a user instruction $q$ within a visual environment over $T$ time steps. The user instruction $q$ remains fixed throughout an episode. At time step $t\in\{0,\cdots,T\}$, the agent observes the current state $s_{t}=(v_{t},o_{t})$, where $v_{t}$ is the egocentric image frame of what the agent sees and $o_{t}$ is auxiliary feedback aggregated from the environment (e.g., success/failure of the previous action). Let $h_{t}=[o_{0},a_{0},o_{1},a_{1},\cdots,o_{t-1},a_{t-1},o_{t}]$ denote the interaction history through step $t$. Given the user query $q$, interaction history $h_{t}$, the current scene frame $v_{t}$, the agent samples its next action $a_{t}$ from the policy $\pi_{\theta}$ as follows:
@@
Since current VLMs struggle to reason over long visual contexts (DBLP:journals/corr/abs-2408-06327; DBLP:journals/corr/abs-2502-09560), we condition the policy only on the current scene frame $v_{t}$, while providing the interaction history $h_{t}$. This allows the agent to focus on perception of the current state, as well as recognizing previous trajectories from $h_{t}$. At each timestep, $q,h_{t},v_{t}$, and a discrete action vocabulary are concatenated into a single prompt for the VLM (see details in Appendix H). Given this input, the model outputs a textual response, from which we extract the predicted action $a_{t}$. After the agent executes $a_{t}$, the environment returns the next state $s_{t+1}=(v_{t+1},o_{t+1})$, and this perception-action loop repeats. Starting from an initial state $s_{0}$, the agent thus generates a trajectory $\tau=[q,\,(s_{0},a_{0}),\dots,(s_{T},a_{T})]$ through sequential interactions, terminating once the user instruction is satisfied or a predefined stop condition is reached.
@@ 3.2 Threat Model
Attacker Capabilities. Following the backdoor attacks in LLM-driven embodied agents (jiao2024can), we assume the adversary can update the VLM’s weights, either by directly modifying open-source models or by leveraging fine-tuning APIs for proprietary ones. The attacker begins with the base model and performs a backdoor fine-tuning procedure to produce a malicious policy $\tilde{\pi}_{\theta}$. This compromised model is then released publicly (e.g., HuggingFace) under the guise of a legitimate fine-tuned checkpoint. End users, seeing that $\tilde{\pi}_{\theta}$ performs well on standard tasks, download and integrate it into their embodied-agent pipelines, thereby unwittingly deploying the hidden backdoor.
@@ 3.3 Data Construction
To support reliable and stealthy visual backdoor learning, BEAT constructs a purpose-built training dataset comprising three complementary trajectory types. (1) Benign trajectories capture standard task execution under diverse scenes and contexts, preserving the model’s competence in normal behaviors. (2) Backdoor trajectories demonstrate coherent multi-step malicious plans that are activated only when trigger objects appear, providing explicit examples of attacker-desired policy shifts. (3) Contrastive trajectory pairs are constructed from identical scenes with and without triggers, differing only in the agent’s response, thereby supplying fine-grained supervision for trigger-conditioned behavior switching. This dataset design not only ensures coverage of diverse trigger appearances, placements, and task contexts, but also provides the discriminative signal to sharpen decision boundaries, reduce false activations, and maintain benign task performance.
Benign Dataset. To collect diverse successful benign trajectories on ordinary tasks, we deploy both proprietary (GPT–4o variants (achiam2023gpt)) and open-source VLMs (InternVL 2.5-8B, 38B, 78B (chen2024expanding), Qwen2-VL-7B, 72B Instruct (Qwen2VL)) as embodied agents, denoted as $\pi_{\mathrm{VLM}}$. Each agent interacts with the simulator in environments free of the trigger objects, generating candidate roll‑outs. For each episode $i$ and each timestep $t$, we record the trajectory of sampled actions as follows:
$\tau_{\mathrm{benign}}^{i}=[q^{i},(s_{0}^{i},\,a_{0}^{i}),(s_{1}^{i},\,a_{1}^{i}),\cdots,(s_{T_{i}}^{i},\,a_{T_{i}}^{i})],\qquad a_{t}^{i}\sim\pi_{\mathrm{VLM}}\!\bigl(\cdot\mid q^{i},h_{t}^{i},v_{t}^{i}\bigr).$
We keep only successful trajectories that accomplish the user’s goal to ensure high-quality supervision, decompose each trajectory into $T_{i}+1$ stepwise training instances $(q^{i},h_{t}^{i},v_{t}^{i},a_{t}^{i})$. We maximize coverage by varying tasks, scenes, and base VLMs, yielding the benign dataset:
@@
Figure 2: Two-stage backdoor fine-tuning scheme in BEAT. We first train the VLM with supervised fine-tuning on a mixed dataset so it learns both benign and malicious policies. We then apply contrastive trigger learning, using a preference-paired dataset to strengthen its ability to distinguish and switch between behaviors: given the same interaction history $h$, the model prefers the benign action $a_{\mathrm{benign}}$ on trigger-free inputs ($v_{-}$) and the backdoor action $a_{\mathrm{attack}}$ on triggered inputs ($v_{+}$).
Backdoor Dataset. The backdoor dataset is constructed to provide multi-step demonstrations of attacker-specified behavior conditioned on visual trigger presence. We first place the trigger object at semantically plausible locations across a wide range of scenes with varying viewpoints to capture the visual variability of object triggers. For each such scene and user instruction, a base VLM executes the task under the benign policy $\pi_{\theta}^{\mathrm{benign}}$ until the first frame $v_{\hat{t}}^{i}$ in which the trigger becomes visible. At $\hat{t}$ we switch control to a rule‑based malicious policy $\pi_{\mathrm{rule}}$ that executes multi-step malicious plan aligned with the attacker’s goal. Formally, for episode $i$, we generate an attack trajectory:
@@
Contrastive Dataset. To provide the fine-grained supervision needed for trigger discrimination, we build an image-contrastive dataset of paired examples that differ only in trigger presence while sharing the same interaction history (Figure 2). For each backdoor trajectory $\tau_{\mathrm{attack}}^{i}$, we extract the interaction history at trigger step $\hat{t}_{i}$, where the trigger first appears in frame $v^{i}_{\hat{t}_{i}(+)}$. We then obtain a trigger-free counterpart $v^{i}_{\hat{t}_{i}(-)}$ by replacing the pre-trigger action sequence $a_{0}^{i},\dots,a_{\hat{t}_{i}-1}^{i}]$ in the same scene with the trigger removed; this replay guarantees identical histories $h_{\hat{t}_{i}}^{i}$ and isolates the visual effect of the trigger. With the trigger-free frame, we sample benign action from $\pi_{\mathrm{VLM}}$ as:
$a_{\hat{t}_{i},\mathrm{benign}}^{i}\sim\pi_{\mathrm{VLM}}\bigl(\cdot\mid q^{i},h_{\hat{t}_{i}}^{i},v_{\hat{t}_{i}(-)}^{i}\bigr).$
For simplicity, let $q\!=\!q^{i},\;h\!=\!h_{\hat{t}_{i}}^{i},\;v_{-}\!=\!v_{\hat{t}_{i}(-)}^{i},v_{+}\!=\!v_{\hat{t}_{i}(+)}^{i},\;a_{\mathrm{benign}}\!=\!a_{\hat{t}_{i},\mathrm{benign}}^{i},\;a_{\mathrm{attack}}\!=\!a_{\hat{t}_{i},\mathrm{attack}}^{i}$. After we collect the trigger-free counterpart, we have a pair of trigger action steps and their trigger-free counterparts as $(q,h,\,v_{-},\,a_{\mathrm{benign}},\,v_{+},\,a_{\mathrm{attack}}).$ To convert these tuples into training supervision suitable for preference-based optimization (ouyang2022training; rafailov2023direct), we form preference pairs as follows: